SHOW WORK

A) 77% of daily drivers use SUV.A sample of 1100 drivers is surveyed and they reported 75% of SUV use. What is the probability of this outcome?

B) An average dancer requires 1.7 liter of water when active. standard deviation of 0.76. A company is planning for a dancing for 50 people and brings 100 liter of water. Probability of the company wont run out of water?

C) amount of water sold daily at a station is uniformly distributed with min of 2,000 gallons, max of 5,000 gallons. probability that the station will sell at least 4k gallons
Answer #1

Solution

(a) Given that n = 1100 , P(SUV drivers), p = 0.77 and sample p = 0.75

Z = (sample p - p)/((p)(1-p)/100)0.5

Z = (0.75-0.77)/((0.77)(0.23)/100)0.5 = -0.4752

probability = 1 - 0.3173 = 0.6827